Question 48
ProbabilityTwo cards are drawn one after the other with replacement from a well-shuffled ordinary pack of 52 cards containing four Aces. Find the probability that they are both Aces.
- A.1/221
- B.1/169
- C.1/52
- D.1/26
- E.1/13
Show the answer and explanation
Answer: B (1/169)
Explanation
With replacement each draw has probability 4/52 = 1/13, so both Aces = 1/13 × 1/13 = 1/169.